A charged particle is launched with a velocity of 5.2x10*m/s at an angle of 35 with respect to a 0.0045T magnetic field. If the magnetic field exerts a force of 0.0026 N on the particle, determine the magnitude of the charge on the particle.
Answer #1

F = q v B sin theta

q = F / (v B sin theta)

= (0.0026) / (5.2 * 104 * 0.0045 * sin 35)

charge on the particle = 1.94 * 10-5 C

= 19.4 muC